Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 16:56 Wednesday 17 March 2021 (16277) Print this report
Comment to Tightening of the wire clamp screws. (16272)

With Hirata-san and Washimi-san.

See pictures in album PR3 Remedying Work.

Summary: after tightening the wire clamps we tested the pitch stability using the Trump method. However, the situation did not change. Pitch still changes a lot when the suspension is shaken.

The work sequence was this:

  1. We tightened the positioning clamp screws for the RM wires to 2 Nm.
  2. We put back the side IRM panels with the OSEMs.
  3. We put the OSEM flags in place.
  4. We released the payload and thew BF side.
  5. We inspected the horizonal OSEM flags visually and saw there was a large negative pitch. H1 and H3 flags were touching the OSEM bodies.
  6. We corrected the pitch with the picomotors and we also modified the roll a little using the readout from V2 and V3 OSEMs, which were within range. The IM became free from touching the IRM.
  7. Transfer function measurements in L, Y and R suggested the IM was free. The coherence was low a low frequencies likely because the IM was still moving a lot due to us working with it.
  8. We set up an oplev at the back of the mirror.
  9. We applied the Trump procedure.

Upon the procedure, according to the oplev, the mirror moved in pitch approximately by -5.8 mrad, whereas, according to OSEMs V2 and V3, the IM moved by -1.04 mrad. See the details of both calculations in this notebook page.

Note: although it was clarified in the commented entry, it's worth writing here that the screws of the locating clamps for the piano wire holding the mirror are not accesible. See this picture

Search Help